in a study of distances traveled by buses before the first major engine filure a sample of 191 buses resulted in mean of 96700 miles and a stndard deviation of 37500 mules.at the 0.05 level of significance test the manufcturer's claim that the mean distance traveled before major engine failure is more than 90000 miles.


1. Claim:

Ho:

Ha:


2. Level of Significance:

Test -statistics:

Tails in Distribution:


3. Reject Ho if:


4. Compute for the value of the test statistics:


5. Make a decision:


6. State the conclusion in terms of the original problem.


1
Expert's answer
2021-04-14T12:47:59-0400

n = 191

"\bar{X} = 96700"

σ = 37500

1. Claim:

Ho: μ ≤ 90000

Ha: μ > 90000

2. Level of Significance: 0.05

Test- statistic:

"P(X>96700) = P(Z> \frac{96700-90000}{37500/ \sqrt{191}} ) \\

= P(Z> 2.469) \\

= 1 -0.9932 \\

= 0.0068"

3. Reject Ho if: P≤0.05

4. Compute for the value of the test statistics.

0.0068<0.05

5. Make a decision: Reject Ho

6. State the conclusion in terms of the original problem: The mean distance traveled before a major engine failure is more than 90,000 miles.